# AI-Based Stroke Rehabilitation Domiciliary Assessment System with ST-GCN Attention

### v1

@@ Abstract
Effective stroke recovery requires continuous rehabilitation integrated with daily living. To support this need, we propose a home-based rehabilitation exercise and feedback system. The system consists of (1) hardware setup with RGB-D camera and wearable sensors to capture Stroke movements, (2) a mobile application for exercise guidance, and (3) an AI server for assessment and feedback. When Stroke user exercises following the application guidance, the system records skeleton sequences, which are then Assessed by the deep learning model, RAST-G@. The model employs a spatio-temporal graph convolutional network (ST-GCN) to extract skeletal features and integrates transformer-based temporal attention to figure out action quality. For system implementation, we constructed the NRC dataset, include 10 upper-limb activities of daily living (ADL) and 5 range-of-motion (ROM) collected from stroke and non-disabled participants, with Score annotations provided by licensed physiotherapists. Results on the KIMORE and NRC datasets show that RAST-G@ improves over baseline in terms of MAD, RMSE, and MAPE. Furthermore, the system provides user feedback that combines patient-centered assessment and monitoring. The results demonstrate that the proposed system offers a scalable approach for quantitative and consistent domiciliary rehabilitation assessment. Our code is available at https://github.com/LimSuH/NRC-rehab
@@ I Introduction
RECENT advancements in Neurology, particularly in motor control and learning, have revealed different mechanisms that induce changes in brain plasticity and behavior over both short- and long-term periods. Physical rehabilitation can be seen as a form of motor learning that occurs under specific conditions[1, 2, 3, 4], and patients with motor impairments, such as those following a stroke, are capable of limited motor learning, although with variations in learning speed and volume. In particular, usage-based and reward-based learning, which are shaped by habitual, repetitive actions and rewards, play a key role in determining long-term brain and behavioral changes in stroke patients after they are discharged and resume daily activities.[5] Therefore, rehabilitation performed at domestic environments is significant in the patient’s recovery. However, without expert, improper rehabilitation can lead to injury and a professional training is essential for evaluations and feedback in order to provide effective rehabilitation. In this study, we propose AI-based Assessment model called RAST-G@ and a outline of rehabilitation monitoring system that can be used within the Domiciliary Environments. This system predicts the user’s rehabilitation movements RGB-D video data and expert assessment score collected during the system development and experiment. Many existing studies provide limited evaluation for topical motion segments. (ex: Range of Angle) But ours takes the sequence of actions as input and learns it with a Spatio-temporal graph based deep learning method (GCN). Since movements are continuous inputs Through time, temporal feature learning is important to evaluate the quality of actions. We enhanced temporal learning results by introducing Temporal Attention, which has proven the performance in deep learning. evaluation Questionnaires from professional therapists contain the quality of complex actions. RAST-G@, the model that end-to-end deep learning-based rehabilitation Assessment is able to assist healthcare providers and contributes to quantitative and consistent assessments of rehabilitation without the need for separate standardization processes. In this paper, We propose an end-to-end deep learning-based rehabilitation Assessment model. This model assists healthcare providers and contributes to quantitative and consistent assessments of rehabilitation without the need for separate standardization processes. It has the potential to enhance the activation of the rehabilitation and the quality of healthcare.
@@ II-A Human Action Representation
One of the most important factors in a deep learning-based framework is the dataset. The higher quality human behavioral representations within dataset, the more meaningful features can be extracted. However, human behavior is abstract high-level data that is in multiple dimensions. Consequently, extensive research has been devoted to identifying the most suitable representation. Traditionally human action representations have been sensor-based.[6, 7] Sensor-based studies have been widely conducted in the rehabilitations. For example, M. Panwar et al.,[8], attached 3-axis accelerometers to classify arm movements including Reach and retrieving, lifting and Rotating arm. in stroke rehabilitation. Similarly, B. Pogorelc et al.,[9] recognized gait patterns of elderly to enable early detection of health problems. Sensor data are advantageous in rehabilitation contexts where continuous monitoring is required, offering benefits in the cost and robustness against environments compared to RGB vision data. On the other hand, vision-based modalities, while computationally heavier and susceptible to privacy concerns and background noise, have gained attention for their convenience in capturing full-body motion.[10] Sensors are dependent on placement and attachment, and reconstructing full-body correlation structures often requires additional procedures. In contrast, vision-based approaches impose less burden on participants and allow for easier equipment installation. To overcome the limitations of both approaches, recent deep learning Method has adopted more processed representations, particularly skeleton-based data[7, 11]. Skeleton represents the human body as a set of landmarks corresponding to joints connected according to anatomical structure. This representation facilitates the modeling of joint interactions through graph, mitigates weaknesses of vision, such as sensitivity to background noises[12], and occlusion and depth ambiguities can be compensated.
@@ II-B Recognition and Assessment
Recognition tasks are designed to categorize input data, ordinarily sequential Data, into predefined discrete labels. Within the broader domain of human action recognition, these tasks aim to identify and differentiate human actions from motion or sensor-derived signals. Unlike simple classification, however, recognition addresses continuous data that may contain inseparable or irrelevant segments. Many studies have extended deep learning networks designed for recognition tasks. A example is the Spatio-Temporal Graph Convolutional Network (ST-GCN)[13], which has proven particularly effective for skeleton. In skeleton graphs, joints are represented as nodes and anatomical connections as edges, enabling effective modeling of human actions. Since human action is characterized by the coordinated movements of multiple joints, ST-GCNs are good for capturing their relational dependencies. Building upon spatio-temporal skeleton recognition[14], demonstrated that even simple data augmentation techniques can yield state-of-the-art performance. Duan et al.,[15] proposed a model that transforms skeletons into heatmaps and trains them using 3D Convolutional Networks. This approach mitigates the impact of outliers arising from pose estimation errors, thereby achieving strong performance in action recognition. When applied to rehabilitation, recognition focuses on categorizing rehabilitation-specific movements. This constitutes a high-dimensional and comprehensive task; as atypical data, it involves not only the identification and segmentation of meaningful rehabilitation-related behaviors from patients’ irregular movements but also their transformation into clinically interpretable information. Assessment is related to recognition, as high-quality evaluation requires clear distinction and the meaningful features. In rehabilitation, assessment tasks aim to measure the quality of movement. Prior implementations have attempted to map performed movements to corresponding quality scores. However, a major challenge in this domain is that rehabilitation assessment metrics are subjective and may significantly depending on the evaluation tool employed. With this issue, earlier research used distance Metrics to produce quantitative assessments. Examples include methods based on Mahalanobis distance[16] and Dynamic Time Warping (DTW) algorithms [17, 18, 19, 20]which measure the distance between a reference movement and a patient’s movement. While these approaches provide objective values, they are sensitive to noise and lack sufficient verification as reliable rehabilitation metrics. M. Capecci et al.,[21] demonstrated that DTW assessments deviate from clinical assessment, highlighting the importance of temporal dynamics in the assessment process. Other attempts include mapping handcrafted features to clinical scores [22] though such approaches require multiple models and incur additional computational cost. The Fugl-Meyer Assessment (FMA) is also widely used as a quantitative clinical Metric. However, its scoring system −{0: The patient cannot perform the task, 1: The patient performs the task partially, 2: The patient fully performs the task.} −is limited in granularity, potentially making item descriptions overly complex and less adaptable for automated assessment.
@@ II-C Deep Learning Network for Human Action
Traditionally, human action data have been modeled in deep learning frameworks using a combine of spatial (skeleton-based) and temporal networks. The temporal dimension is particularly critical, as the sequential nature of movements often has a more direct impact on model performance than static body configurations. Liao et al.,[23] simultaneously incorporates spatial structure and temporal dynamics. Their method progressively learned five human body subsets and integrated them to capture overall motion, ensuring that temporal dependencies were reflected in the model. Because of limited rehabilitation datasets, they also employed a temporal pyramid structure that subsampled motion sequences. However, the convolutional networks used for spatial learning were not well suited for skeleton data, where body joints are deeply interrelated.[24] This limitation naturally motivated the adoption of Graph Convolutional Networks (GCNs), which can preserve connectivity of the human body.[25] Deb et al.,[26] presented a notable example introducing GCNs into rehabilitation. Learning sequential data into a Spatio-Temporal GCN (ST-GCN), A single model could evaluate multiple types of movements. From prior works, their also incorporated Long Short-Term Memory (LSTM) networks for temporal modeling, inspired by Liao et al. As a variant of recurrent neural networks (RNNs), it designed for the vanishing gradient and long-term memory problems by gating mechanisms. Nonetheless, LSTMs still suffer from limitations in modeling very long-term dependencies[27] and their complexity leads to overfitting and interpretability issues.[28]
@@ III Methods
the objective of this study is to learn a regression function $f_{\theta}:X\to Y$ that maps an input motion sequence $X\in\mathbb{R}^{N\times C\times T\times V\times M}$ to a continuous assessment score $Y$, as provided by clinical experts. Here, $N$ is the batch size, $C$ is the channel dimension, $T$ is the sequence length, $V$ is the number of keypoints (25 joints in Fig. 1(a)), and $M$ is the number of instances. The output $Y$ is a scalar score annotated by clinicians for both stroke patients and non-disabled participants. Based on this formulation, we describe the acquisition of clinically relevant rehabilitation data and the design of assessment items, and the proposed network that combines ST-GCN and Transformer-based attention modules to effectively capture the spatio-temporal characteristics of skeleton sequences.
@@ III-A Collecting Dataset
Domiciliary rehabilitation systems demand both efficiency and intuitive usability. To achieve this, we designed the system pipeline such that each module operates in coordination. Because rehabilitation data for stroke patients is costly and difficult to acquire, whenever a user performs rehabilitation exercises, the software simultaneously conducts recognition and assessment while recording data for training. The overall system is illustrated in Fig. 2. The hardware setup includes a mobile tablet, an RGB-D camera, and IMU sensors. The software provides exercise guidance, records the corresponding data, and presents evaluation results. The recording function manages exercise execution and data collecting, while the assessment function connects with the server to present assessment score.
@@ III-A1 Rehabilitation Actions Data
Using The proposed system, We collected a multimodal rehabilitation dataset. The hardware setup included an Intel RealSense D435i depth camera and Movella Xsens Dot sensors. From these, RGB-D videos were captured, and acceleration·Gyroscope $(x,y,z)$ was collected from attached devices to both wrists. For the Assessment model presented in this paper, only RGB-D data was used, in order to maintain compactness and interpretability. The Classes consisted of 15 movements in Table. I. Including both Activities of Daily Living (ADL) and Range of Motion (ROM) tasks, chosen to evaluate upper-limb functional ability.[29] The movement set was confirmed under consultation with faculty members in physical therapy. As in Table. II., data was collected from two participant groups: non-disabled (ND) and stroke patients. Skeleton keypoints were estimated from the Captured RGB-D videos. We adopted 25 keypoints configuration, which is the most widely used representation in skeleton-based human motion analysis. This ensures compatibility with existing skeleton-based benchmarks such as OpenPose [30] and COCO WholeBody[31], both of which are based on 25 full-body keypoints. Since many prior works have leveraged these benchmarks for human motion analysis, the 25 keypoints not only enables performance comparison but also aligns with well-studied graph construction methods. This experiment for collecting dataset was pproved by the IRB of Gachon university before proceeding (IRB no. 1044396-202306-HR-102-02)
@@ 5 MovingCan Moving a tool UNI Can
6 Writing Writing with one hand while holding paper with the other BIA Pencil
@@ III-A2 Assessment Score
The most significant component of the assessment dataset was evaluation items. As noted in prior studies [17, 18, 19, 20, 21, 22] and in the representative rehabilitation dataset UI-PRMD [32], Most existing approaches reference the movements of ND as ideal motions. However, according to Birke et al.,[33], The main goal of neurological rehabilitation is the improvement of function and independence. In other words, rehabilitation assessment focuses on achieving functional activities and participation[34], rather than the mere removal of impairments. Consequently, the quality of a top grades movement performed by a Stroke cannot be equated with that of a ND. This underscores the need for stroke-centered Assessment criteria. Although detailed clinical assessments can be applied, subjectivity remains a concern. Therefore, there is a need to establish new set of evaluation items with upper-limb rehabilitation tasks. The KIMORE[35] provides an illustrative example of balancing quantitative and clinical evaluation: it employs 10 items, each scored on a 5-point Likert scale, resulting in a maximum score of 50 per trial. Table. II. summarizes the questionnaire items used in NRC(ours).We developed evaluation items specifically designed for the 15 upper-limb movements. The items were initially drafted through discussions among three licensed physiotherapists with more than five years of clinical experience, and validated by two Ph.D. members in physical therapy. Item 2 and 3 were under the assumption that each movement is performed only once, while 8,9 and 10 address seated tasks involving upper-limb usage. In particular, item 8 evaluates fine motor skills such as handling small objects, which are essential for assessing activities of daily living (ADL).
@@ III-B Model Structure
As reported by Deb et al.,[26], Graph Convolutional Networks (GCNs) are effective as a framework for assessing rehabilitation movements. Human motion can be represented as a graph, with skeleton keypoints as nodes and anatomical connections as edges. However, this corresponds only to the spatial representation of the human skeleton. human movements are expressed as sequential information, a temporal dimension is required in the network. For this reason, spatio-temporal GCNs (ST-GCNs) have been shown to be effective in human action deep learning tasks[13]. Rehabilitation movement assessment can be regarded as a subtask of human action and thus inherits these properties; in our methods, through skeleton-based ST-GCN, spatial skeleton information is connected along the temporal dimension. The model structure is illustrated in Fig. 3. The skeleton consists of 25 keypoints, including the upper limbs and both hands. The first ST-GCN block takes as input the skeleton coordinates $X\in\mathbb{R}^{N\times C\times T\times V\times M}$ and the graph $G=(V,E,A)$ where adjacency matrix is $A\in\mathbb{R}^{K\times V\times V}$, where $C$ denotes the channel dimension, $T$ the frame-sequence length, and $V$ the number of keypoints. The skeleton graph is constructed for each frame from human joint coordinates, and convolution is performed across both spatial and temporal dimensions. For hop distance $k$, the adjacency matrix $A$ is normalized as $\tilde{A}^{(l)}_{k}=D_{k}^{-1/2}\bigl(A_{k}\odot E_{k}^{(l)}+I\bigr)D_{k}^{-1/2}$. At the $l$th layer, each ST-GCN unit is defined as in  (1) and (2). Equation (3) representates an ST-GCN Block.
@@
Figure 3: Overview of RAST-G@ Model Structure. (a) Skeleton sequence and Score input, (b) Outputs 3d feature map $X^{\prime}\in\mathbb{R}^{({N\times V})\times T\times C}$ via STGCN Block from skeleton spatio-temporal graph. (c) Structure of a single ST-GCN unit. After GCN and TCN layers, followed by Batch Normalization and ReLU; the output is concatenated with the initial input through residual connections. (c) Transformer attention block. From the ST-GCN output, the query (Q), key (K), and value (V) are obtained, and the attention scores are computed as in (4). (d) The prediction $Y$ is compared with the ground-truth scores, $\hat{Y}$
Fig. 3(c) illustrates a single ST-GCN unit. All units, except for the first, operate with residual structure. After each convolution along the spatial and temporal axes, Batch Normalization and ReLU activation are applied. the output is concatenated with the initial input to preserve the original information. While ST-GCN is powerful in extracting local and sequential patterns, it has the limitation that all joints and temporal segments are treated with equal weight. In actual human actions or rehabilitation movements, certain joints (e.g., wrist, shoulder) or time points (e.g., the beginning or end of a movement) are often more important. Transformer has demonstrated the effectiveness of self-attention mechanisms in deep learning[36]. We adopt Transformer-based network within RAST-G@ to enable score assessment as illustrated in Fig. 3(d). Multi-head attention is applied across the temporal axis for each joint. This still allows temporal features to be learned, but unlike LSTM-based models that aggregate global temporal information across joints, it performs joint-specific temporal attention. Since this does not require additional adjacency matrix operations, it reduces computational overhead while remaining effective for long-term memory. To perform attention along the temporal axis, the Temporal Attention;(TA) module reshapes the input $X$ to $X^{\prime}\in\mathbb{R}^{({N\times V})\times T\times C}$ For head $h$, the query, key, and value matrices are computed as $Q_{h}$$X^{\prime}W^{Q}_{h}$, $K_{h}$$X^{\prime}W^{K}_{h}$, and $V_{h}$$X^{\prime}W^{V}_{h}$. The output is computed as a weighted sum of the values, where the weight assigned to each value is computed by a compatibility function of the query with the corresponding key. The input consists of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Then Compute the dot products of the query with all keys, divide each by $\sqrt{d_{k}}$ and apply a softmax function to obtain the weights on the values.[37]
@@
After learning the attention weights for all sequences, the model outputs a single scalar score representing the quality of the movement. regression head is employed to predict scores over a wide range and diverse cases. The regression module consists of an average pooling layer followed by a linear layer. Pooling aggregates spatio-temporal features into a compact representation $z^{\prime}\in\mathbb{R}^{N\times C\times 1\times 2}$, and the linear layer gets this feature to a scalar score $Y$, in (4). The final model output $Y$ is compared with the ground-truth scores assigned by expert physiotherapists. For model optimization, the Huber Loss, (5) is used with delta($\delta$)=0.1 allowing a selective application of MSE and MAE. When the error is below MSE emphasizes sensitivity to small deviations, while for errors exceeding MAE mitigates the influence of outliers.

### v2

@@ Abstract
Effective stroke recovery requires continuous rehabilitation integrated with daily living. To support this need, we propose a home-based rehabilitation exercise and feedback system. The system consists of (1) hardware setup with RGB-D camera and wearable sensors to capture stroke movements, (2) a mobile application for exercise guidance, and (3) an AI server for assessment and feedback. When a stroke user exercises following the application guidance, the system records skeleton sequences, which are then assessed by the deep learning model, RAST-G@ (Rehabilitation Assessment Spatio-Temporal Graph ATtention). The model employs a spatio-temporal graph convolutional network to extract skeletal features and integrates transformer-based temporal attention to figure out action quality. For system implementation, we constructed the NRC dataset, include 10 upper-limb activities of daily living (ADL) and 5 range-of-motion (ROM) collected from stroke and non-disabled participants, with Score annotations provided by licensed physiotherapists. Results on the KIMORE and NRC datasets show that RAST-G@ improves over baseline in terms of MAD, RMSE, and MAPE. Furthermore, the system provides user feedback that combines patient-centered assessment and monitoring. The results demonstrate that the proposed system offers a scalable approach for quantitative and consistent domiciliary rehabilitation assessment. Our code is available at https://github.com/LimSuH/NRC-rehab
@@ I Introduction
RECENT advancements in neurology, particularly in motor control and learning, have revealed different mechanisms that induce changes in brain plasticity and behavior over both short- and long-term periods. Physical rehabilitation can be seen as a form of motor learning that occurs under specific conditions[20, 15, 29, 24], and patients with motor impairments, such as those following a stroke, are capable of limited motor learning, although with variations in learning speed and volume. In particular, usage-based and reward-based learning, which are shaped by habitual, repetitive actions and rewards, play a key role in determining long-term brain and behavioral changes in stroke patients after they are discharged and resume daily activities.[13] Therefore, rehabilitation performed at domestic environments is significant in the patient’s recovery. However, without expert, improper rehabilitation can lead to injury and a professional training is essential for evaluations and feedback in order to provide effective rehabilitation. In this study, we propose AI-based assessment model called RAST-G@ and a outline of rehabilitation monitoring system that can be used within the domiciliary environments. This system predicts the user’s rehabilitation movements RGB-D video data and expert assessment score collected during the system development and experiment. Many existing studies provide limited evaluation for topical motion segments such as range of angle. But ours takes the sequence of actions as input and learns it with a spatio-temporal graph based deep learning method. Since movements are continuous inputs through time, temporal feature learning is important to evaluate the quality of actions. We enhanced temporal learning results by introducing temporal attention, which has proven the performance in deep learning. Evaluation questionnaires from professional therapists contain the quality of complex actions. RAST-G@, the model that end-to-end deep learning-based rehabilitation assessment is able to assist healthcare providers and contributes to quantitative and consistent assessments of rehabilitation without the need for separate standardization processes. In this paper, we propose an end-to-end deep learning-based rehabilitation assessment model. This model assists healthcare providers and contributes to quantitative and consistent assessments of rehabilitation without the need for separate standardization processes. It has the potential to enhance the activation of the rehabilitation and the quality of healthcare.
@@ II-A Human Action Representation
A high-quality dataset enable the extraction of more informative features. However, human behavior is abstract, which has motivated substantial research on effective data representations. Early work primarily relied on sensor-based modalities.[7, 48] Sensor-based approaches are widely used in rehabilitation because they support continuous monitoring and are often cost-effective and robust to environmental variation when compared with RGB video. For instance, Panwar et al.[33], attached tri-axial accelerometers to classify upper-limb movements, including reaching and retrieving, lifting, and arm rotation in stroke rehabilitation. Pogorelc et al.[34] similarly analyzed wearable sensor signals to identify gait patterns in older adults to facilitate early detection of health issues.
Despite these advantages, sensor-based methods depend on correct device placement and attachment. Modeling whole-body coordination may require additional processing or instrumentation. Vision-based modalities offer an alternative with simple setup. However, They often require more computation and can be raise privacy problems. They can also be sensitive to background noise.[37] To balance these trade-offs, recent deep learning studies increasingly adopt structured, processed representations, particularly skeleton-based data.[48, 43]. Skeleton data represent the body as joint landmarks connected according to anatomical topology. This representation naturally supports graph-based modeling of joint interactions. It can also reduce sensitivity to background noise[8], relative to RGB-based approaches, while partially mitigating occlusion and depth ambiguity through the use of kinematic structure and temporal context.
@@ II-B Recognition and Assessment
Recognition tasks categorize sequential input data, to predefined discrete labels. In human action recognition, the goal is to identify and distinguish actions from motion trajectories or sensor-derived signals. In contrast to classification tasks, recognition commonly operates on continuous streams that may include irrelevant segments or transitions, which necessitates temporal features and robustness to ambiguity. A broad range of deep learning architectures has been developed for action recognition. Among them, the Spatio-Temporal Graph Convolutional Network called ST-GCN[46] has shown strong performance on skeleton-based data. In a skeleton graph, joints are modeled as nodes and anatomical connections are represented as edges, which enables explicit modeling of coordinated joint motion. Because human actions are defined by structured interactions among joints, ST-GCN is suited to capturing relational dependencies. Building on spatio-temporal skeleton recognition, relative work has also reported that even simple augmentation strategies can yield state-of-the-art performance.[44] Duan et al.[12] proposed an alternative representation that converts skeleton sequences into heatmaps and trains a 3D convolutional network. This approach reduces outliers from pose estimation errors and improves action recognition performance.
In rehabilitation, recognition aims to classify rehabilitation-specific movements. This is challenging because patient motion is often irregular, requiring identification and segmentation of meaningful behaviors and conversion into clinically interpretable outputs. The assessment task is related to recognition because accurate evaluation requires clear discrimination of movements and extraction of meaningful features. In rehabilitation, assessment aims to quantify movement quality. However, rehabilitation metrics are often subjective and can vary substantially across evaluation tools. To address this limitation, earlier studies adopted distance-based metrics to derive quantitative assessments. Representative examples include approaches based on Mahalanobis distance[14] and Dynamic Time Warping called DTW[39, 52, 1, 35], which compute dissimilarity between a reference movement and a patient’s movement. Although these methods yield objective values, they can be sensitive to noise and have limited validation as clinically reliable metrics. Capecci et al.[5]. reported that DTW-based scores can diverge from clinician scores.
Other work has mapped handcrafted features to clinical scores[23], but such pipelines often require multiple components and increase computational cost. The Fugl–Meyer Assessment (FMA) is a widely adopted traditional rehabilitation instrument. Because it provides quantitative scores, it can also be considered a practical and digitization-friendly target metric for training AI-based assessment models. However, its three-level scoring scheme−{0: unable to perform, 1: partial performance, 2: full performance}−provides limited granularity. This scale requires detailed descriptions of motion and evaluation criteria, making it less suitable for automated assessment.
@@ II-C Deep Learning Network for Human Action
Traditionally, deep learning frameworks for human action have modeled motion by combining spatial encoders with temporal modeling modules. The temporal dimension is particularly important because the sequential ordering and dynamics of movement frequently influence performance more strongly than static pose configurations. Liao et al.[27] proposed a framework that jointly incorporates spatial structure and temporal dynamics. Their method progressively learned features from five body-part subsets and then integrated them to represent whole-body motion, thereby encouraging the model to capture temporal dependencies. Given the limited size of available rehabilitation datasets, they additionally employed a temporal pyramid strategy that subsamples sequences to improve learning efficiency and robustness. However, conventional convolutional networks used for spatial feature extraction are not inherently matched to skeleton data, where joints interdependence through anatomical connectivity[11]. This limitation has motivated the adoption of Graph Convolutional Networks called GCN, which preserve the anatomical connectivity of the human body during model learning[30]. Building on this direction, Deb et al.[10] introduced GCN-based modeling into rehabilitation assessment. By learning sequential skeleton data with a ST-GCN, their approach enabled a single model to evaluate multiple action classes.
Following earlier work, they also incorporated Long Short-Term Memory(LSTM) networks for temporal features, drawing inspiration from Liao et al.[19] LSTMs are a variant of Recurrent Neural Networks(RNN) that use gating mechanisms to mitigate vanishing gradients and to retain information over longer horizons. Nevertheless, LSTM can still be limited in capturing very long-range dependencies, and their architectural complexity may increase the risk of overfitting and reduce interpretability.[21]
@@ III Methods
The objective of this study is to learn a regression function $f_{\theta}:X\to Y$ that maps an input motion sequence $X\in\mathbb{R}^{N\times C\times T\times V\times M}$ to an assessment score $Y$, which is a scalar value provided by clinicians for both stroke patients and Non-Disabled (ND) participants. Here, $N$ is the batch size, $C$ is the channel dimension, $T$ is the sequence length, $M$ is the number of instances, and $V$ is the number of keypoints shown in Fig. 1(a).
The output $Y$ is a scalar score annotated by clinicians for both stroke and ND participants. Based on this formulation, we describe the acquisition of clinically relevant rehabilitation data and the design of assessment items, and the proposed deep learning network that combines ST-GCN blocks and transformer-based Temporal Attention modules to effectively capture the spatio-temporal characteristics of skeleton sequences.
@@ III-A Collecting Dataset
Domiciliary rehabilitation systems require both computational efficiency and intuitive usability. To meet these requirements, we designed an end-to-end pipeline in which all modules operate in a coordinated manner. Because stroke rehabilitation data are costly and difficult to obtain, the software performs recognition and assessment whenever a user completes an exercise, while also recording the resulting data for model training. The overall system architecture is illustrated in Fig. 2. The hardware configuration consists of a mobile tablet, an RGB-D camera, and IMU sensors. The software comprises a recording module and an assessment module. The recording module manages exercise execution and data acquisition, whereas the assessment module communicates with a deep learning assessment model on the server to compute and display the predicted score.
@@ III-A1 Rehabilitation Actions Data
Using the proposed system, we collected a multimodal rehabilitation dataset. The hardware setup included an Intel RealSense D435i depth camera and Movella Xsens Dot sensors. We captured RGB-D videos and collected accelerometer and gyroscope signals $(x,y,z)$ from sensors attached to both wrists. For the assessment model presented in this paper, we used only RGB-D data to maintain a compact and interpretable pipeline. As summarized in Table I, data were collected from two participant groups: ND participants and stroke patients. Table II lists the 15 action classes included in the dataset, comprising Activities of Daily Living (ADL) and Range Of Motion (ROM) tasks selected to evaluate upper-limb functional ability.[31] The movement set was finalized in consultation with faculty members in physical therapy. For the model input, skeleton keypoints were estimated from the captured RGB-D videos. We adopted a 25-keypoint configuration, which is widely used in skeleton-based human analysis.[18] This configuration supports compatibility with commonly used skeleton-based representations and benchmarks, including OpenPose[3] and COCO WholeBody[45]. It also facilitates direct performance comparisons with prior work and aligns with well-established graph construction strategies. This data collection study was approved by the IRB of Gachon University.(IRB no. 1044396-202306-HR-102-02)
@@ 5 MovingCan Moving a tool UNI Can
6 Writing Writing with one hand while holding paper with the other hand BIA Pencil
@@ III-A2 Assessment Score
The definition of evaluation items is a central component of an assessment dataset. As noted in prior studies[39, 52, 1, 35, 5, 23] and in the representative rehabilitation dataset UI-PRMD[41], many existing approaches treat movements performed by ND as the reference for ideal motion. However, according to Birke et al.[2], the primary goal of neurological rehabilitation is to improve function and independence. Rehabilitation assessment therefore emphasizes the achievement of functional activities and participation[25], rather than the mere reduction of impairments. Consequently, even a high-scoring movement performed by a stroke patient should not be directly equated with that of an ND participant, which motivates the need for stroke-specific assessment criteria. Although detailed clinical assessments can be adopted, subjectivity remains a concern. This underscores the necessity of a new set of evaluation items for upper-limb rehabilitation tasks. The KIMORE[6][4] provides an illustrative example of balancing quantitative scoring with clinical interpretation. It employs 10 items rated on a 5-point Likert scale, yielding a maximum score of 50 per trial. Table III summarizes the questionnaire items used in our NRC dataset. We developed evaluation items specifically for the 15 upper-limb movements in our dataset. The items were initially drafted through discussions among three licensed physiotherapists with more than five years of clinical experience and were subsequently validated by two Ph.D researchers in physical therapy. Items 2 and 3 assume that each movement is performed only once, whereas Items 8, 9, and 10 address seated tasks involving upper-limb use. In particular, item 8 evaluates fine motor skills, such as handling small objects, which are essential for ADL.
@@ III-B Model Structure
Deb et al.[10] reported that GCN provide an effective framework for assessing rehabilitation exercises. Human body can be formulated as a graph, where skeleton keypoints are treated as nodes and anatomical connections as edges. However, this formulation primarily captures the spatial structure. Because human movements are inherently sequential, the network must also model temporal dynamics. For this reason, ST-GCN have shown strong performance in human action modeling tasks[46]. Rehabilitation movement assessment can be viewed as a subtask of human action analysis and thus benefits from modeling methods, including GCN and ST-GCN. In our approach, a skeleton-based ST-GCN links spatial joint information across time to capture the spatio-temporal characteristics of rehabilitation actions. The overall model architecture is illustrated in Fig. 3. The skeleton representation consists of 25 keypoints, including upper-limb and hand joints. The first ST-GCN block takes as input the skeleton coordinates $X\in\mathbb{R}^{N\times C\times T\times V\times M}$ and the graph $G=(V,E)$ with adjacency matrix $A\in\mathbb{R}^{K\times V\times V}$. Here, $C$ denotes the channel dimension, $T$ the sequence length, and $V$ the number of keypoints. The skeleton graph is constructed for each frame from the joint coordinates, and convolution is performed across both spatial and temporal dimensions. For hop distance $k$, the adjacency matrix $A$ is normalized as $\tilde{A}^{(l)}_{k}=D_{k}^{-1/2}\bigl(A_{k}\odot E_{k}^{(l)}+I\bigr)D_{k}^{-1/2}$. At the $l$th layer, each ST-GCN unit is defined as in  (1) and (2). Equation (3) representates an ST-GCN Block.
@@
Figure 3: Overview of RAST-G@ model structure. (a) Skeleton sequence and Score input, (b) Outputs 3d feature map $X^{\prime}\in\mathbb{R}^{({N\times V})\times T\times C}$ via ST-GCN block from skeleton spatio-temporal graph. (c) Structure of a single ST-GCN unit. After GCN and TCN layers, followed by batch normalization and ReLU; the output is concatenated with the initial input through residual connections. (d) Temporal Attention(TA) block. From the ST-GCN output, the query (Q), key (K), and value (V) are obtained, and the attention scores are computed as in (4). The prediction $Y$ is compared with the ground-truth scores, $\hat{Y}$.
Fig. 3(c) illustrates a single ST-GCN unit. All units, except for the first, employ residual connections. After each convolution along the spatial and temporal axes, Batch Normalization and ReLU activation are applied. The output is then combined with the initial input to preserve the original information. While ST-GCN is effective at extracting local spatio-temporal patterns, it treats all joints and temporal segments as equally important. In practical human actions and rehabilitation motions, however, certain joints (e.g., the wrist and shoulder) and specific time points (e.g., the onset or offset of a movement) often contribute more substantially to the overall motion.
Transformers have demonstrated the effectiveness of self-attention mechanisms in deep learning[9]. We incorporate a transformer-based module for temporal feature learning within RAST-G@ to predict assessment scores, as illustrated in Fig. 3(d). Multi-head attention is applied along the temporal axis independently for each joint. This design enables feature learning while performing joint-specific temporal attention, in contrast to LSTM-based models that aggregate temporal information across joints. Because this module does not require additional adjacency-matrix operations, it reduces computational overhead while remaining effective for long-term defendencies. To perform attention along the temporal axis, the Temporal Attention (TA) module reshapes the input $X$ to $X^{\prime}\in\mathbb{R}^{({N\times V})\times T\times C}$ For head $h$, the query, key, and value matrices are computed as $Q_{h}$$X^{\prime}W^{Q}_{h}$, $K_{h}$$X^{\prime}W^{K}_{h}$, and $V_{h}$$X^{\prime}W^{V}_{h}$. The output is computed as a weighted sum of the values, where the weight assigned to each value is determined by a compatibility function between the query and the corresponding key. The input consists of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Next, the dot products of the query with all keys are computed, each is scaled by $\sqrt{d_{k}}$, and a softmax function is applied to obtain attention weights over the values.[42]
@@
After learning attention weights over the input sequence, the model outputs a single scalar score representing action quality. A regression head is employed to predict scores across a wide range of conditions. The regression module consists of an average pooling layer followed by a linear layer. The pooling operation aggregates spatio-temporal features into a compact representation $z^{\prime}\in\mathbb{R}^{N\times C\times 1\times 2}$, and the linear layer maps this representation to a scalar score $Y$, as in (4). The final model output $Y$ is compared with the ground-truth scores assigned by expert physiotherapists. For model optimization, the Huber Loss, (5) is used with delta($\delta$)=0.1, which combines the properties of MSE and MAE. When the prediction error is smaller than $\delta$, the loss behaves like MSE and is sensitive to small deviations. When the error exceeds $\delta$, it behaves like MAE, thereby reducing the influence of outliers.